OpenStudy (anonymous):

If f(x) = ∣(x2 − 6)(x2 + 2)∣, how many numbers in the interval 1 ≤ x ≤ 2 satisfy the conclusion of the mean value theorem?

OpenStudy (anonymous):

well first thing's first its really |(x^2-6)(X^2+2)|

OpenStudy (anonymous):

also in that interval given how many numbers can apply to the mean value theorem which the formula for it is

OpenStudy (anonymous):

you want to prove the mean value theorem saying that \(\large\color{black}{ \tt f'(c)=\frac{f(a)-f(b)}{a-b} }\). In other words, to porve that the derivative at x=c, is the slope of the secant between a and b, over the interval, \(\large\color{black}{ \tt [a,b] }\) So I think you would need to know that it is continuous at this interval \(\large\color{black}{ \tt [1,2] }\). Well, all I got so far, is that it is a polynomial, and the absolute value is defined And polynomial is contious on \(\large\color{black}{ \tt (-\infty,\infty) }\)by definiton.

OpenStudy (anonymous):

Right. And I thought plugging in the intervals in the equation would solve something but im not positive.

OpenStudy (anonymous):

But before anything I take the derivative of both polynomials. Should i do product rule?

OpenStudy (anonymous):

\(\Large\color{black}{ \tt f'(c)= \frac{f(2)-f(1)}{2-1} }\) is the slope between the endpoints.

OpenStudy (anonymous):

I haven't actually done derivatives of an absolute value though.

OpenStudy (anonymous):

I would just distribute and do the power rule.

OpenStudy (anonymous):

well i just suppose that it will be a positive number, but absolute values has a cusp, so wouldn't that make the equation undefined at a certain point?

OpenStudy (anonymous):

did you distribute it already?

OpenStudy (anonymous):

Yes I did, now the undefined values. Yes, lets see. and \(\Large\color{black}{ \tt x^4-4x^2-12=0 }\) \(\Large\color{black}{ \tt x^4-4x^2+4=16 }\) Let, \(\Large\color{black}{ \tt x^2=a }\) \(\Large\color{black}{ \tt a^2-4a+4=16 }\) \(\Large\color{black}{ \tt (a-2)=\pm 4}\) \(\Large\color{black}{ \tt a=\pm 4+2}\) \(\Large\color{black}{ \tt a=6,~~-2}\)

OpenStudy (anonymous):

\(\Large\color{black}{ \tt x=\pm\sqrt{6},~~~\pm i\sqrt{2}}\)

OpenStudy (anonymous):

and the sqrt(6) is greater than 2

OpenStudy (anonymous):

where did you get \[x^4-4x^2+4=16\]?

OpenStudy (anonymous):

I added 16 to both sides.

OpenStudy (anonymous):

where did the 16 come from sorry i got lost.

OpenStudy (anonymous):

I wanted to make the left side into a perfect square polynomial.

OpenStudy (anonymous):

(which I did)

OpenStudy (anonymous):

okay got it

OpenStudy (anonymous):

but, you want to set the derivative equal to the slope of the secant.

OpenStudy (anonymous):

tell me what is: f(2)=? f(1)=?

OpenStudy (anonymous):

f(2)=12 f(1)=15

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

By the way the answer choices I have if it is not to late is Zero One Two Three

OpenStudy (anonymous):

connection snapped, but f(2) and f(1) are correct.

OpenStudy (anonymous):

Alright great so far.

OpenStudy (anonymous):

\(\LARGE\color{black}{ \tt f'(c) = \frac{ f(2) - f(1) }{ 2 - 1 } }\) plug in your values. And here you can use this, copy the gray and use it.... `\(\LARGE\color{black}{ \tt f'(c) = \frac{ f(2) - f(1) }{ 2 - 1 } }\)`

OpenStudy (anonymous):

(\tt s just the font)

OpenStudy (anonymous):

so let's see

OpenStudy (anonymous):

fill in your values for f(2) and f(1)... and simplify.

OpenStudy (anonymous):

\[12-15/2-1\]

OpenStudy (anonymous):

Use my thing in gray copy it, and go for it...

OpenStudy (anonymous):

and i'd get \[-3\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So, now, set the derivative equal to that.

OpenStudy (anonymous):

okay so what did we say the derivative was?

OpenStudy (anonymous):

I was thrown off by it.

OpenStudy (anonymous):

\(\LARGE\color{black}{ \tt f'(x)=\frac{(x^4-4x^2-12)(4x^3-8x)}{|4x^2-4x^2-12|} }\)

OpenStudy (anonymous):

\(\LARGE\color{black}{ \tt -3=\frac{(x^4-4x^2-12)(4x^3-8x)}{|4x^2-4x^2-12|} }\)

OpenStudy (anonymous):

solve for x....

OpenStudy (anonymous):

i got 84

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

How do you get only 1 solution?

OpenStudy (anonymous):

wasn't i supposed to plug it in?

OpenStudy (anonymous):

I have a question, @freckles , can we cancel the \(\LARGE\color{black}{ \tt 4x^2-4x^2-12 }\) with the \(\LARGE\color{black}{ \tt |4x^2-4x^2-12| }\) ?

OpenStudy (anonymous):

I mean since we are only dealing with positive xs? and in this case, not in general.

OpenStudy (freckles):

if 4x^2-4x^2-12>0 then |4x^2-4x^2-12|=4x^2-4x^2-12 if 4x^2-4x^2-12<0 then |4x^2-4x^2-12|=-(4x^2-4x^2-12) so it would depend on the restrictions of x and if that restriction is a subset of the first inequality or the second inequality I mentioned

OpenStudy (anonymous):

\(\LARGE\color{black}{ \tt f'(x)=\frac{(x^4-4x^2-12)(4x^3-8x)}{|4x^4-4x^2-12|} }\) Correction.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x%5E4-4x%5E2-12%3E0 x>sqrt(6) x<-sqrt(6) So in this case, where our interval is 2 and 1,. we would need to multiply the top x^4-4x^2-12 times -1, and then cancel it with the absolute value. I mean like \(\LARGE\color{black}{ \tt f'(x)=\frac{(x^4-4x^2-12)(4x^3-8x)}{|4x^4-4x^2-12|} }\) \(\LARGE\color{black}{ \tt f'(x)=\frac{-(x^4-4x^2-12)(8x-4x^3)}{|4x^4-4x^2-12|} }\) \(\LARGE\color{black}{ \tt f'(x)=\frac{\bcancel{\cancel{\color{red}{-(x^4-4x^2-12)}}}(8x-4x^3)}{\bcancel{\cancel{\color{red}{|4x^4-4x^2-12|}}} } }\)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

But how would the x^4 and 4x^4 cancel?

OpenStudy (freckles):

If f(x) = ∣(x2 − 6)(x2 + 2)∣, how many numbers in the interval 1 ≤ x ≤ 2 satisfy the conclusion of the mean value theorem? f(x)=(x^2+2)|x^2-6| x^2+2>0 for all real x f is not differentiable at sqrt(6) or -sqrt(6) g(x)=x^2-6 looks like:  it is negative between -sqrt(6) and +sqrt(6) so that is where fbi got the - sign so since our function is restricted on 1 to 2 we have f(x)=-(x^2+2)(x^2-6) differentiate that by product rule or multiply it all out then differentiate f(x)=-(x^4-4x^2-12) f'(x)=-4x^3+8x on the interval (1,2)